# MEt3R: Measuring Multi-View Consistency in Generated Images

### v1

@@ Abstract
Figure 1: We introduce MEt3R, a metric for multi-view consistency between pairs of generated images, which is independent of image quality, image content, and does not require camera poses. Left: generated images from different generative models, conditioned on the first frame, with MEt3R score map indicating levels of inconsistencies between consecutive images $i$ and $i+1$. Right: pair-wise consistency scores, evaluated for consecutive frames in a sliding window, averaged over multiple sequences. The pattern in MV-LDM’s consistency clearly shows artifacts from using anchor frames that are generated first, highlighting the high signal-to-noise ratio of MEt3R.
@@ 1 Introduction
Generative models, such as diffusion [13, 33] or flow-based [20] models, are trained to sample from a given data distribution, which makes them ideal candidates for stochastic inverse problems, such as reconstruction from incomplete information [43, 10, 36]. However, they raise the inherent challenge that for individual samples no ground truth is available to measure the quality of generations with pair-wise distance metrics. As a result, metrics such as FID [12], KID [1] and CMMD [15] have been proposed to measure the quality of generated images without the need of a paired ground truth.
@@
Recently, a trend is to repurpose video [14, 3] and image [34, 28] diffusion models for generation of 3D scenes and objects, by generating multiple views from different camera poses [10, 50, 39], with or without given images as conditioning. In comparison to direct generation of 3D representations [24, 6, 30], such multi-view generative models can be trained on images and videos and their pixel-aligned representation allows for more efficient models and better scalability. However, they have only a weak to non-existent inductive bias to produce actually 3D consistent results, which is of large importance for the subsequent lift into 3D. A reliable metric to evaluate the multi-view consistency of such generations is desperately needed to advance these models further. Luckily, similar to general image quality, 3D consistency between views can be evaluated without the existence of paired ground truth data. Existing metrics [46] though, fail to reliably perform such evaluation, as shown in Fig. 2. In this work, we propose a metric to measure 3D consistency, which is independent of the specific scene and model used to generate the images, works under changing lighting conditions, does not require camera poses, is differentiable, and is a gradual measure of consistency instead of a binary one.
MEt3R utilizes DUSt3R [40] to obtain dense reconstructions from image pairs in a common 3D space. It then projects features of one image into the view of the other using the reconstructed point maps and computes feature similarity between the obtained images. As feature extractor, DINO [4] + FeatUp [9] are used to obtain features that are more robust to view-dependent effects, such as lighting, and which can be compared quantitatively. We further introduce an open-source multi-view latent diffusion model (MV-LDM) to be used in our studies, which is able to generate high-quality and consistent scenes. MEt3R is evaluated in different scenarios to validate its usefulness and robustness. It is used to benchmark existing methods that generate multi-view images of scenes, with and without intermediate 3D representation, as well as our MV-LDM. We can show that MV-LDM performs well in the quality vs. consistency trade-off and find that MEt3R is a reliable metric that aligns well with the expectations of measuring consistency. In contrast to previous metrics, it can distinguish perfectly consistent from almost consistent sequences and can capture fine-grained changes in consistency over time.
@@
  • the first metric for measuring multi-view consistency of generated views without given camera poses,
  • a comprehensive analysis of existing methods that generate multi-view images of scenes and videos, and
@@ 3 MEt3R: Measuring Consistency
In this section, we introduce MEt3R, our feed-forward metric to measure multi-view consistency. Given two images as input, a metric for multi-view consistency should (1) penalize image pairs that are not consistent, and (2) must not penalize pairs that are consistent but deviate from a given ground truth or do not follow a desired distribution. Thus, we develop MEt3R to be orthogonal to image quality metrics, e.g., FID [12], and to pixel-wise reconstruction metrics, e.g. PSNR.
@@
Given the scores $S(\mathbf{I}_{1},\mathbf{I}_{2})$ and $S(\mathbf{I}_{2},\mathbf{I}_{1})$, we can define MEt3R as
[…]
@@ 3.1 Stereo Reconstruction with DUSt3R
The core of our method relies on pose-free stereo reconstruction of pixel-aligned point clouds. Given image pair $\mathbf{I}_{1}$, $\mathbf{I}_{2}$, the DUSt3R [40] model $\Psi$ regresses pixel-aligned 3D point clouds $\mathbf{X}_{1}\in\mathbb{R}^{H\times W\times 3}$ and $\mathbf{X}_{2}\in\mathbb{R}^{H\times W\times 3}$:
@@ 3.2 High-Resolution Feature Similarity
Concretely, we first use DINO [4] to obtain semantic features for $\mathbf{I}_{1}$ and $\mathbf{I}_{2}$. Then, since the corresponding feature maps are of low resolution and do not represent detailed structures, we upsample them using FeatUp [9], which employs an image-adaptive upsampling i.e., a stack of Joint Bilateral Upsamplers (JBUs) that learned to upsample low resolution feature maps from DINO. It uses the high resolution image to transfer high frequency information to the upsampling process, allowing the upsampled features to faithfully reconstruct and preserve important details.
Let $\mathbf{F}_{1}$ and $\mathbf{F}_{2}$ denote the upsampled DINO features from images $\mathbf{I}_{1}$ and $\mathbf{I}_{2}$, respectively. Then, we unproject both features into 3D space using the DUSt3R point maps and subsequently reproject them onto the camera frame of $\mathbf{I}_{1}$:
@@ 4 Multi-View Latent Diffusion Model
Additionally to our metric, we provide an open-source multi-view latent diffusion model (MV-LDM). It is inspired by the architecture of CAT3D [10], which is not publicly available. While CAT3D is trained on top of proprietary image/video diffusion models, we initialize our model with StableDiffusion and train it on the openly available dataset RealEstate10k [49]. For a detailed description of MV-LDM, we refer to the appendix Sec. A. Our code and model are publicly available for further research.
@@
MV-LDM encodes images into a latent space using a pre-trained VAE encoder from StableDiffusion 2.1  [28]. Then, camera ray encodings are concatenated to the latent images, providing camera pose information before being processed by the diffusion model. We take a pre-trained StableDiffusion UNet model, add attention between views in the latent space, and finetune it on RealEstate10k videos for 2M iterations. MV-LDM works with a total of 5 views at a time consisting of $N$ conditioning and $M$ target views.
@@
We adopt the anchored generation strategy from CAT3D [10]. When generating many views of a scene, the generation process starts with sampling 4 anchor images for widely distributed cameras, conditioned on a single input image. Then, in the second step, the remaining views are generated and conditioned on the closest anchor images along with the initial input image. The goal of the anchoring strategy is to prevent accumulating errors that often occur when generating target views autoregressively, conditioned on the previously generated views. When generating with anchors, the accumulation of errors can be effectively limited. We analyze the effect on consistency and image quality in Sec. 5.3.
Figure 4: Metric comparison. We compare MEt3R against TSED, SED, and FVD by computing average per-frame (/-segment for FVD) scores over a large number of generated sequences. MEt3R is able to capture nuanced differences in consistency of DFM, MV-LDM, and real videos, while TSED rates them all very similar. Unlike MEt3R, SED does not capture increasing inconsistency for PhotoNVS and DFM. MEt3R is able to capture the influence of anchor views in MV-LDM (c.f. Sec. 4 and appendix Sec. A) as structured high-frequency patterns. For MEt3R, the standard deviation gradually increases, starting from a small value, which is expected behavior due to conditioning on the first frame and is not the case for the other metrics.
@@ 5 Experiments
    How consistent are the outputs of existing generative models for multi-view and video generation?
@@ 5.1 Experimental Setup
To evaluate MEt3R, we consider two sets of baselines for multi-view and video generation models. In addition, we categorize the multi-view generation methods into three general classes: 1) single-view, 2) multi-view, and 3) 3D diffusion models.
Multi-view Generation Models
We consider GenWarp [31], which is a single view image-to-image inpainting diffusion model, and PhotoNVS [46], which is an autoregressive multi-view generation model that generates a single view at a time conditioned on the previous. Moreover, we consider DFM [36], which is a 3D diffusion method that incorporates a neural radiance field into the architecture of an image diffusion model, forcing the rendered novel views to be 3D consistent by design. Finally, MV-LDM, our own open-source multi-view diffusion model, coupled with cross-view attention (c.f. Sec. 4), generates multiple novel views at a time, resulting in a stronger 3D prior than single-view methods, i.e., GenWarp and PhotoNVS. We refer to appendix Sec. C for further details on these baselines.
@@
Video Generation Models
We take Stable Video Diffusion (SVD) [2], Ruyi-Mini-7B [35] and I2VGen-XL [47], which are standard open source video diffusion models that can generate videos with a single input image with an additional text prompt.
Dataset
To faithfully benchmark with MEt3R, we collect 100 image sequences from the RealEstate10K [49] test set. We take the first image for each sequence as the initial input, followed by 80 target poses, which the multi-view generation models generate. We perform consecutive pairwise evaluations on the generated images in a sliding-window fashion. In this way, we: 1) allow maximal projection area and more overlapping pixels to evaluate; 2) cover regions that are extrapolated and not visible in the input image; and 3) investigate the evolution of pairwise consistency as the camera pair moves further away from the input image. We set a standard resolution of $256^{2}$ as input to MEt3R. In case of DFM [36], we upsample from $128^{2}$, and for GenWarp[31], we downsample from $512^{2}$ bilinearly. Similarly, we use identical test sequences for video diffusion models but limit the number of generated frames to 48 due to memory restrictions. Note that we do not have explicit camera control over the generation and, therefore, are not equivalent in camera trajectories. In addition, the generated videos differ in resolution and aspect ratios, which we resize accordingly to the closest resolution of $256^{2}$ while maintaining the aspect ratio.
@@ 5.2 Validating MEt3R
We validate the efficacy of MEt3R by computing the lower bound that the baseline MEt3R must follow. Intuitively, we can evaluate MEt3R on a dataset of real video sequences to obtain the lower bound. Although real video sequences are assumed to be perfectly 3D consistent, a lower bound slightly above zero is observed, attributed to errors in point map alignment from DUSt3R and small 3D inconsistencies in DINO features. The results for real videos are shown together with the results from the multi-view generation baselines in Fig. 4.
Methods MEt3R $\downarrow$ TSED $\uparrow$ SED $\downarrow$ FVD $\downarrow$ FID $\downarrow$ KID $\downarrow$
GenWarp  [31] 0.120 0.674 1.398 1312.7 29.80 0.0033
PhotoNVS [46] 0.069 0.996 0.479 1498.7 43.67 0.0081
MV-LDM (Ours) 0.036 0.998 0.405 945.8 37.29 0.0064
DFM [36] 0.026 0.990 0.346 1174.6 73.02 0.0324
I2VGen-XL [47] 0.051 - - 1722.6 66.88 0.0161
Ruyi-Mini-7B [35] 0.047 - - 850.5 42.67 0.0071
SVD [2] 0.033 - - 674.6 48.33 0.0111
Table 1: Quantitative comparison. Average MEt3R alongside FID [12], KID [1], FVD [38], SED [46] and TSED [46]. (a) Plot comparing MEt3R with FID and FVD. (b) Quantitative comparison of multi-view and video generation baselines. Among multi-view methods, DFM achieves the best consistency in MEt3R and SED but the worst in FID and KID due to their sensitivity to blur artifacts, aligning with the visual impression in Figs. 51519. While GenWarp delivers the best image quality, it has the worst consistency. In contrast, our MV-LDM achieves a favorable position in the image quality vs. consistency trade-off for multi-view generation. Unlike TSED and SED, MEt3R applies to generated video as it does not require camera poses. Though a weak correlation between MEt3R and FVD is observed in video generation, this does not extend to multi-view generation. As FVD assesses the overall quality and temporal coherency across several videos, MEt3R evaluates pairwise 3D consistency within individual video sequences without relying on a ground truth dataset.
@@
We compare MEt3R with existing metrics to measure 3D consistency. As baselines, we consider SED [46], TSED [46], and FVD [37] for multi-view generation. In Fig. 4, we plot per image-pair scores for all generated frames, averaged over 100 sequences. For FVD, we compare the distributions of image segments by splitting the sequences into chunks of 10 frames each. We find that MEt3R, SED, and FVD increase as we progress through the image-pair sequence, suggesting a decrease in consistency, which is also qualitatively visible in Fig. 5. Although TSED captures this trend for GenWarp [31], it does not report a meaningful separation for other baselines. Unlike TSED and SED, MEt3R captures the gradual decrease in consistency for PhotoNVS [46] and MV-LDM. For GenWarp, MEt3R captures this trend more accurately, starting with a lower score and standard deviation, as the first frame provides stronger conditioning for closer views with a larger overlap, resulting in better consistency. Furthermore, we observe sudden periodic spikes for MV-LDM in MEt3R and SED, attributed to transition artifacts when we switch between anchors during sampling (c.f. Sec. 4 and A.2). Unlike all other metrics, DFM [36] is consistently worse than MV-LDM in terms of FVD. Since DFM is supposed to be 3D consistent by design (c.f. Sec. 5.1 and C), this shows that FVD is sensitive to blurry samples and does not evaluate multi-view consistency. Ideally, a larger sample size is preferred to accurately capture and compare the underlying distribution of generated and ground-truth image sequences [37], to which FVD is sensitive. Therefore, it cannot be applied at the level of individual image pairs.

### v2

@@ Abstract
Figure 1: We introduce MEt3R, a metric for multi-view consistency between pairs of generated images, which is independent of image quality and content and does not require camera poses. Left: generated images from different generative models, conditioned on the first frame, with MEt3R score map (Cons. Error) indicating levels of inconsistencies between consecutive images $i$ and $i+1$. Right: pair-wise consistency scores, evaluated for consecutive frames in a sliding window, averaged over multiple sequences. The pattern in MV-LDM’s consistency clearly shows artifacts from using anchor frames that are generated first, highlighting the high signal-to-noise ratio of MEt3R.
@@ 1 Introduction
Generative models, such as diffusion [14, 34] or flow-based [21] models, are trained to sample from a given data distribution, which makes them ideal candidates for stochastic inverse problems, such as reconstruction from incomplete information [45, 11, 38]. However, they raise the inherent challenge that, for individual samples, no ground truth is available to measure the quality of generations with pairwise distance metrics. As a result, metrics such as FID [13], KID [1], and CMMD [17] have been proposed to measure the quality of generated images without the need for a paired ground truth.
@@
Recently, a trend is to repurpose video [15, 3] and image [35, 28] diffusion models for generation of 3D scenes and objects, by generating multiple views from different camera poses [11, 52, 41], with or without given images as conditioning. Compared to direct generation of 3D representations [26, 6, 31], such multi-view generative models can be trained on images and videos, and their pixel-aligned representation allows for more efficient models and better scalability. However, they have only a weak to non-existent inductive bias to produce actually 3D consistent results, which is of large importance for the subsequent lift into 3D. A reliable metric to evaluate the multi-view consistency of such generations is critically needed to advance these models further. Fortunately, similar to general image quality, 3D consistency between views can be evaluated without the existence of paired ground-truth data. Existing metrics such as TSED [48], though, fail to reliably perform such evaluation, as shown in Fig. 2. In this work, we propose a metric to measure 3D consistency, which is independent of the specific scene and model used to generate the images, works under changing lighting conditions, does not require camera poses, is differentiable, and is a gradual measure of consistency instead of a binary one.
MEt3R utilizes DUSt3R [42] to obtain dense reconstructions from image pairs in a common 3D space. It then projects features of one image into the view of the other using the reconstructed point maps and computes feature similarity between the obtained images. As feature extractors, DINO [4] + FeatUp [10] are used to obtain high-resolution features from input images that are robust view-dependent effects, such as lighting, while preserving semantics and image-level structures, to quantify 3D consistency. We further introduce an open-source multi-view latent diffusion model (MV-LDM) to be used in our studies, which is able to generate good quality and consistent scenes. MEt3R is evaluated in different scenarios to validate its usefulness and robustness. It is used to benchmark existing methods that generate videos and multiple views of objects and scenes, with and without an intermediate 3D representation, as well as our MV-LDM. We show that MV-LDM performs well in the quality vs. consistency trade-off and find that MEt3R is a reliable metric that aligns well with the theoretical expectations of consistency among the different classes of scene generation methods. In contrast to previous metrics, it can distinguish perfectly consistent from almost consistent sequences and can robustly capture fine-grained changes in consistency over time.
@@
  • a simple yet effective metric for measuring 3D consistency of generated views without given camera poses,
  • a comprehensive analysis of existing methods that generate videos and multiple views of objects and scenes, and
@@ 3 MEt3R: Measuring Consistency
In this section, we introduce MEt3R, our feed-forward metric to measure multi-view consistency. Given two images as input, a metric for multi-view consistency should (1) penalize image pairs that are not consistent, and (2) must not penalize pairs that are consistent but deviate from a given ground truth or do not follow a desired distribution. Thus, we develop MEt3R to be orthogonal to image quality metrics, e.g., FID [13], and to pixel-wise reconstruction metrics, e.g., PSNR.
@@
Given the scores $S(\mathbf{I}_{1},\mathbf{I}_{2})$ and $S(\mathbf{I}_{2},\mathbf{I}_{1})$, we can define MEt3R as,
$\textnormal{MEt3R}(\mathbf{I}_{1},\mathbf{I}_{2})=1-\frac{1}{2}\Big(S(\mathbf{I}_{1},\mathbf{I}_{2})+S(\mathbf{I}_{2},\mathbf{I}_{1})\Big)\textnormal{,}$ (1)
which gives $\textnormal{MEt3R}(\cdot,\cdot)\in[0,2]$, lower is better, due to $S(\cdot,\cdot)\in[-1,1]$, and is symmetric. We found $S$ to already behave approximately symmetric. Thus, in practice, $\textnormal{MEt3R}(\cdot,\cdot)$ can also be approximated well by only computing one direction of $S$ in case of runtime constraints. We now provide the details for the DUSt3R reconstruction in Sec. 3.1 and feature similarity in Sec. 3.2
@@ 3.1 Stereo Reconstruction with DUSt3R
The core of our method relies on pose-free stereo reconstruction of pixel-aligned point clouds. Given an image pair $\mathbf{I}_{1}$, $\mathbf{I}_{2}$, the DUSt3R [42] model $\Psi$ regresses pixel-aligned 3D point clouds $\mathbf{X}_{1}\in\mathbb{R}^{H\times W\times 3}$ and $\mathbf{X}_{2}\in\mathbb{R}^{H\times W\times 3}$:
@@ 3.2 High-Resolution Feature Similarity
Concretely, we first use DINO [4] to obtain semantic features for $\mathbf{I}_{1}$ and $\mathbf{I}_{2}$. Then, since the corresponding feature maps are of low resolution and do not represent detailed structures, we upsample them using FeatUp [10], which employs an image-adaptive upsampling, i.e., a stack of Joint Bilateral Upsamplers (JBUs) that learned to upsample low-resolution feature maps from DINO. It uses the high resolution image to transfer high frequency information to the upsampling process, allowing the upsampled features to faithfully reconstruct and preserve important details.
Let $\mathbf{F}_{1}$ and $\mathbf{F}_{2}$ denote the upsampled DINO features from images $\mathbf{I}_{1}$ and $\mathbf{I}_{2}$, respectively. We unproject both features into 3D space using the DUSt3R point maps and subsequently reproject them onto the camera frame of $\mathbf{I}_{1}$:
@@ 4 Multi-View Latent Diffusion Model
Additionally to our metric, we provide an open-source multi-view latent diffusion model (MV-LDM). It is inspired by the architecture of CAT3D [11], which is not publicly available. While CAT3D is trained on top of proprietary image/video diffusion models, we initialize our model with Stable Diffusion 2.1 [28]. For a detailed description of MV-LDM, we refer to the appendix Sec. A. Our code and model are publicly available for further research.
@@
MV-LDM encodes images into a latent space using a pre-trained VAE encoder from Stable Diffusion. Then, ray maps are concatenated to the input latents, providing camera pose information. We take the 2D UNet architecture, add attention between views at each UNet block, and finetune the full model on RealEstate10k for 1.65M iterations.
@@
We adopt the anchored generation strategy from CAT3D. When generating many views of a scene, the process starts by sampling four anchor images for widely distributed cameras, conditioned on a single input image. Then, in the second step, the remaining views are generated and conditioned on the closest anchor and the initial input image. The goal of the anchoring strategy is to prevent accumulating errors that often occur when generating target views autoregressively, conditioned on the previously generated views. When generating with anchors, the accumulation of errors can be effectively limited. We analyze the effect on consistency and image quality in Sec. 5.3.
Figure 4: Metric comparison. We compare MEt3R against TSED, SED, FWS (PSNR), and FVD by computing average per-frame (/-segment for FVD) scores over many generated sequences. MEt3R can capture nuanced differences in the consistency of DFM, MV-LDM, and real videos, while TSED rates them all very similarly. Unlike MEt3R, SED does not capture increasing inconsistency for PhotoNVS and DFM. MEt3R also captures the influence of anchor views in MV-LDM (c.f. Sec. 4 and A.2) as structured high-frequency patterns. For MEt3R, the standard deviation gradually increases, starting from a small value. This behavior is expected due to the proximity of initial frames to the conditioning frame (c.f. Sec. 5.3.3) and is not the case for the other metrics.
@@ 5 Experiments
    How consistent are the outputs of existing generative models for multi-view and video generation of objects and scenes?
@@ 5.1 Experimental Setup
To evaluate MEt3R, we consider three sets of baselines for multi-view, video, and object-level generation models. In addition, we categorize the multi-view generation methods into three general classes: 1) single-view, 2) multi-view, and 3) 3D diffusion models.
Multi-view Generation Models.
We consider GenWarp [32], which is a single-view inpainting diffusion model, and PhotoNVS [48], which is a two-view generation model that generates a single view at a time conditioned on the previous. Moreover, we consider DFM [38], which is a 3D diffusion method, and MV-LDM, our own open-source multi-view latent diffusion model, coupled with cross-view attention (c.f. Sec. 4 and A). For more details on baselines, we refer to Sec. D in the appendix.
@@
Video Generation Models.
We take Stable Video Diffusion (SVD) [2], Ruyi-Mini-7B [36] and I2VGen-XL [49], which are standard open source video diffusion models that can generate a full video from a single input image.
Object-Level Generation Models.
From object-level methods, we compare EpiDiff [16], SyncDreamer [22], and VideoMV [53]. EpiDiff and SyncDreamer employ an underlying multi-view diffusion model, while VideoMV uses a video diffusion model to generate novel views of objects.
Dataset.
To faithfully benchmark with MEt3R, we collect 100 image sequences from the RealEstate10K [51] test set. We take the first image for each sequence as the initial input, followed by 80 target poses, which the multi-view generation models generate. We perform consecutive pairwise evaluations on the generated images in a sliding-window fashion. In this way, we: 1) allow maximal projection area and more overlapping pixels to evaluate; 2) cover regions that are extrapolated and not visible in the input image; and 3) investigate the evolution of pairwise consistency as the camera pair moves further away from the input image. We set a standard resolution of $256^{2}$ as input to MEt3R. In the case of DFM, we upsample from $128^{2}$, and for GenWarp, we downsample from $512^{2}$ bilinearly. Similarly, we use identical test sequences for video diffusion models but limit them to 48 frames to address memory constraints. Note that we do not have explicit camera control over the generation and, therefore, are not equivalent in camera trajectories. The generated videos also differ in resolution, which we resize accordingly to the closest resolution of $256^{2}$ while preserving the aspect ratio. For object-level methods, we use Google Scanned Objects (GSO) [9] dataset, which consists of $360^{\circ}$ views of objects. We subdivide the range $0^{\circ},360^{\circ}]$ into 16 frames at $256^{2}$ resolution, forming a closed loop with the frame at $0^{\circ}$ as the conditioning. Both EpiDiff and SyncDreamer generate 16 frames, while VideoMV generates a fixed set of 32 frames, which we uniformly downsample to 16.
@@ 5.2 Validating MEt3R
We validate the efficacy of MEt3R by computing the lower bound that the baselines must follow. Intuitively, we can evaluate MEt3R on a dataset of real video sequences. Although they are assumed to be perfectly 3D consistent, a lower bound slightly above zero is observed, attributed to errors in point map alignment from DUSt3R [42] and small 3D inconsistencies in DINO [4] features. The results for both multi-view generation baselines and real videos are shown in Fig. 4.
Methods MEt3R $\downarrow$ TSED $\uparrow$ SED $\downarrow$ FVD $\downarrow$ FID $\downarrow$ FWS (PSNR) $\uparrow$
GenWarp  [32] 0.120 0.674 1.398 1312.7 29.80 21.41
PhotoNVS [48] 0.069 0.996 0.479 1498.7 43.67 25.10
MV-LDM (Ours) 0.036 0.998 0.405 945.8 37.29 28.46
DFM [38] 0.026 0.990 0.346 1174.6 73.02 39.56
I2VGen-XL [49] 0.050 - - 1722.6 66.88 28.62
Ruyi-Mini-7B [36] 0.047 - - 850.5 42.67 28.01
SVD [2] 0.032 - - 674.6 48.33 29.93
Table 1: Quantitative comparison. Average MEt3R alongside TSED [48], SED [48], FVD [40], FID [13], and FWS (PSNR). (a) Plot comparing MEt3R with FID and FVD. (b) Quantitative comparison of multi-view and video generation baselines. Among multi-view methods, DFM achieves the best consistency in MEt3R, FWS (PSNR), and SED but the worst in FID. We attribute the low FID and high PSNR to blurry renderings, both of which are sensitive, whereas GenWarp delivers the best image quality with worse consistency but with a lower signal-to-noise ratio. In contrast, our MV-LDM achieves a favorable position in the image quality vs. consistency trade-off for multi-view generation. Unlike TSED and SED, MEt3R applies to generated video as it does not require camera poses.
@@
We compare MEt3R with existing metrics to measure 3D consistency. As baselines, we consider SED [48], TSED [48], and FVD [39] for multi-view generation. We also compare with several variants of flow warping score (FWS) using RAFT [37] to warp one frame to another and compute PSNR, SSIM, LPIPS, and RMSE, among which we show the results for PSNR in Fig. 4. For the remaining, we refer to Sec. B in the appendix. In Fig. 4, we plot per image-pair scores for all generated frames, averaged over 100 sequences. For FVD, we compare the distributions of image segments by splitting the sequences into chunks of 10 frames each. We find that MEt3R, SED, FWS (PSNR), and FVD increase as we progress through the image-pair sequence, suggesting a decrease in consistency, which is qualitatively visible in Fig. 5. Although TSED captures this trend for GenWarp [32], it does not report a meaningful separation for other baselines. Unlike TSED and SED, MEt3R captures the gradual decrease in consistency for PhotoNVS [48] and MV-LDM. For GenWarp, MEt3R captures this trend more accurately, starting with a lower score and standard deviation, as the first frame provides stronger conditioning for closer views with a larger overlap, resulting in better consistency. Furthermore, we observe sudden periodic spikes for MV-LDM in MEt3R, FWS (PSNR), and SED, attributed to transition artifacts when we switch between anchors during sampling (c.f. Sec. 4 and A.2). Unlike all other metrics, FVD cannot be applied to image pairs and requires a collection of frames. Ideally, a larger sample size is preferred to accurately capture and compare the underlying distribution of the generated and ground-truth image sequences [39], to which FVD is sensitive. Moreover, both FVD and FWS (PSNR) are sensitive to blur. Specifically, DFM achieves worse FVD, which is supposed to be 3D consistent by design (c.f. Sec. 5.1 and D), while the real video, which is perfectly 3D consistent, gets worse than DFM in PSNR, as shown in Fig. 4.